### v5

@@ 4 Lero Research Ireland centre for Software Research and University of Limerick, Ireland
Classification or Prompting: A Case Study on Legal Requirements Traceability
@@ 1 Introduction
Technological advancements are significantly transforming software development across diverse domains, such as healthcare Caruana:15 . Software applications and automated assistants have become integral to our daily lives Zhan:22 . This evolution, driven by recent breakthroughs in artificial intelligence (AI), has led to increasing complexity in software systems ahmad2023requirements ; Feldt:18 . As technology progresses, regulations are adapting in parallel to ensure that software systems are developed in line with ethical and legal standards. For example, the general data protection regulation (GDPR) GDPR is enforced since 2018 to address concerns about privacy and data protection. Despite being introduced by the European Union (EU), the GDPR has a global effect, impacting organizations (and software) outside the EU as long as they handle personal data of EU residents.
@@
Requirements Engineering (RE) plays a pivotal role in this landscape. RE is concerned with specifying and maintaining software requirements that outline the properties and functions of a system-to-be Pohl:11 . Legal compliance of software systems against applicable provisions can be addressed at different stages of software development. One scenario is to explicitly identify legal requirements early during the requirements elicitation phase, answering the question: “What legal obligations need to be satisfied by the system for it to be compliant?”. The elicited legal requirements can then be integrated into the software development process, while maintaining trace links to the source legal provisions. As an alternative scenario, requirements engineers may need to verify the compliance of existing software systems against legal provisions in a post-deployment stage, as new regulations have become applicable. In this case, they must answer the question “Does the system satisfy the regulation?”. To do so, engineers must analyze the regulation, identify applicable legal provisions, and then trace software requirements to these statements. Both alternatives rely on requirements traceability analysis, an essential RE activity concerned with the identification and maintenance of trace links between requirements and other artifacts within the software development lifecycle Meyer:22 . legal requirements traceability (LRT) is a special case where requirements are traced to provisions in a regulation and is the focus of this paper.
@@
To illustrate this concept, consider the following example. Imagine a fictional mobility app named WeMobilize, which helps users book and share cab rides. Originally a non-EU startup, WeMobilize is expanding to the EU and hence must comply with the GDPR. This example is particularly relevant as many businesses are globalizing and must adapt to data protection laws in different jurisdictions. Fig. 1 shows how WeMobilize’s requirements (labeled REQ1 – REQ5) can be traced to data protection policies in the GDPR GDPR . We identify trace links to provisions in GDPR for REQ1 and REQ3 – REQ5, visualized as dashed lines in black. REQ2 has no trace link to GDPR in our example since it does not involve processing users’ personal data.
@@
REQ1 involves collecting user’s personal information and must therefore be traced to two provisions, namely REG_DIR (related to the direct collection of personal information) and REG_CON (related to the explicit soliciting of users’ consent). Currently, consent is not part of REQ1, which prevents identifying a trace link with REG_CON—a missing trace link is visualized with a red dashed line in the figure. Failing to identify this trace link entails a possible breach of GDPR. Therefore, deploying WeMobilize as-is, without accounting for provisions in GDPR, can lead to potential reputational and financial losses caused by violating the GDPR. LRT can help identify potential non-compliance issues at early stages but requires not only legal expertise but also substantial manual effort. Developing automated support is therefore beneficial to assist engineers and analysts in identifying applicable trace links.
@@
Requirements traceability is a well-explored problem in the RE literature, e.g., wang2018requirements ; tufail2017systematic . However, the extensive research on requirements traceability is not directly applicable to LRT due to the challenges outlined above. Despite the serious consequences of non-compliance, LRT has received limited attention from the community. Cleland-Huang et al. cleland:2010 ; Gibiec:2010 proposed a classifier that predicts trace links by computing the likelihood of requirements being traced to provisions based on indicator terms found in both provisions and requirements. Guo et al. Guo:17 focused on bridging the terminology gap between provisions and software requirements. They examined three methods, including the one by Cleland-Huang et al. mentioned above, and two others based on web-mining and ontologies. The proposed methods aim to expand the terminology of the provisions by adding additional terms to better identify trace links.
@@
Existing traceability approaches for the LRT task, mentioned above, have several limitations. First, they leverage a statistical model based on TF-IDF (Term Frequency-Inverse Document Frequency), which primarily relies on lexical overlap between documents, where similarity is computed from the presence and frequency of overlapping words. This approach is mostly effective when the source and target texts share similar vocabulary. Still, they struggle to capture deeper semantic relationships, which are required in LRT, as the language used in legal texts often differs significantly from that in technical requirements. Second, existing techniques have yet to fully leverage the recent advances in natural language processing (NLP), evolving from latent semantic methods such as Latent Semantic Indexing (LSI) deerwester1990indexing and Latent Dirichlet Allocation (LDA) blei2003latent , which uncover hidden semantic structures through concept and topic modeling, toward more contextually rich approaches. We posit that (large) language models (LMs or LLMs) and contextual sentence embedding architectures offer capabilities beyond lexical or shallow syntactic matching, enabling them to capture nuanced legal and technical semantics crucial to LRT. While some recent work has applied language models to traceability tasks lin2021traceability , these efforts have not been tailored to the distinctive demands of aligning with legal requirements—such as interpreting domain-specific terminology, resolving cross-references, or bridging differences in stylistic and structural conventions between legal and technical texts. Architectures like Sentence Transformers can generate semantically rich representations that are better aligned with these needs but remain underutilized in this context. Third, prompt engineering has not been given sufficient attention; existing work often relies on overly simple prompts ronanki2024requirements ; hey2025requirements ; ge2025cross or on querying multiple LLMs fuchss2025beyond , which may not always be feasible in real-world deployments due to computational and cost constraints. Fourth, the evaluation is based on a single benchmark that may not fully reflect the complexity of the legal domain in practice. To address these limitations, we propose two novel approaches in this paper, based on recent NLP technologies that utilize the Transformer architecture Vaswani:17 and LLMs. Similar to existing work, both approaches aim to predict trace links, and we assess their performance on a realistic scenario beyond the commonly used benchmark dataset.
@@
(3) To further confirm its performance, we test $\mathcal{K}$ashif on new unseen requirements documents covering diverse domains and requirements types. These requirements are traced to the GDPR, a more complex regulation with 26 provisions pertaining to personal data protection that must be adhered to in software requirements. On this dataset, the base ST model (without additional fine-tuning) yields an average recall of 15%. In comparison, a pre-trained sentence transformer, with no exposure to the requirements traceability task, yields a nearly zero recall, as elaborated in Section 4.7. Driven by this observation, we propose our second solution, the final contribution of this paper, as explained next.
@@ 2 Background
Language Models (LMs). Language Modeling in NLP involves computationally determining the probability distribution of word sequences Jurafsky:20 . Given a sequence of words, an LM predicts the most likely next word, enabling it to generate text Alexandrescu:06 . For example, an LM would predict “Mat” as the most likely next word in the input sequence, “The cat sits on the [WORD]”. LMs are trained on large corpora of texts to estimate these probability distributions accurately. State-of-the-art LMs are based on transformer architecture, which leverages self-attention mechanisms to weigh the significance of different parts of an input text relative to a given position Vaswani:17 . The attention mechanism determines which words in a sentence are more important based on the context and gives them more “attention”. For instance, in the sentence “Mary, who used to live in Paris, loves wine.”, the attention is on Mary and wine. Building on transformer architectures, the Sentence Transformers framework (ST) Reimers:19 offers a set of pre-trained models designed to encode longer text sequences, such as sentences or paragraphs, into dense vector representations within a high-dimensional space. They produce contextual embeddings that capture the overall semantic essence of an entire input sequence.
More recently, generative LLMs have emerged as transformer-based language models that are scaled up significantly in model size and the amount of training data. Examples on LLMs include OpenAI’s GPT (Generative Pre-trained Transformer) Radford:18 and LLaMa touvron2023llama ; touvron2023llama2 . These models can perform new tasks based on textual instructions (prompts) NEURIPS2020_1457c0d6 .
Machine learning (ML). Supervised learning is one of the most prominent paradigms in ML. In this paradigm, the ML algorithm is provided with labeled training data where each data point consists of an input vector (features) and the corresponding output label (or value). The algorithm learns patterns within the input features to make predictions based on this training. When trained on a sufficiently large dataset, the algorithm refines its predictions to classify the provided labels more accurately. The example ML classification algorithms include random forest, decision tree, support vector machine, and feed-forward neural networks Goodfellow:16 .
@@ 3.2 $\mathcal{K}$ashif
Fig. 2 provides a comprehensive overview of the two phases comprising our approach. Phase A covers steps 1-3 and offers a developer’s perspective, focusing on building a traceability model for solving LRT. Step 1 prepares a training dataset of manually identified trace links. Step 2 selects a pre-trained model to customize for addressing LRT. Step 3 involves fine-tuning the LRT model. Phase B covers steps 4-6 and provides the perspective of an end user (e.g., a requirements analyst) assuming the availability of an LRT model. Step 4 preprocesses the input requirements document (RD). Step 5 applies the LRT model to compute the semantic similarities between each requirement in the RD and each provision. Step 6 predicts trace links. We explain these steps in detail next.
@@
In step 4, we preprocess the input requirements using a simple NLP pipeline composed of two modules, namely Tokenization and sentence splitting. The goal is to decompose the text into separate sentences. In our work, a requirement $r_{i}$ corresponds to a sentence generated by the NLP pipeline, which may or may not be grammatically correct. Using $\mathcal{K}$ashif to solve LRT for multi-sentence requirements is straightforward. A provision is traced to the requirement if it is traced to any sentence thereof. The intermediary output of this step is a set of $n$ requirements ($\mathcal{R}=\{r_{1},r_{2},\ldots,r_{n}\}$) from the input RD.
@@
(a) Constant Threshold: To predict a trace link, we utilize a pre-defined constant threshold, $\theta=0.5$. Specifically, a trace link is predicted if the similarity score exceeds $0.5$. This threshold is considered a reasonable rule of thumb, as evidenced by its previous application in the literature Yao2014 ; Corley2005 . Moreover, this threshold was chosen based on the fine-tuning approach used in $\mathcal{K}$ashif. During training, positive pairs are labeled with a similarity score of 1, while negative pairs are labeled with 0. The model is thus optimized to produce similarity scores closer to 1 for true links and closer to 0 for unrelated pairs. As a result, a threshold of 0.5 naturally serves as a midpoint between these two extremes, making it a natural choice for distinguishing between positive and negative links.
@@
(b) Dynamic Threshold: Another practical method to adjust $\theta$ involves curating a set of negative training examples, i.e., requirements that do not have trace links. These requirements can be sourced from publicly available datasets or from different projects. However, for more accurate results, it is ideal to use requirements from the same project under analysis. Inspired by similarity-based classification proposed in the literature Amaral:21 , we select $\theta$ using the following procedure. For each provision $c_{j}\in\mathcal{C}$, we identify a set of negative training examples ($TR_{j}^{-}$), i.e., requirements $\{r_{1}^{\prime},\ldots,r_{k}^{\prime}\}$ that do not have trace links to $c_{j}$. We then compute the similarity between $r_{i}$ and $TR_{j}^{-}$ and set $\theta$ to the average cosine similarity between $r_{i}$ and $TR_{j}^{-}$. If the similarity between $r_{i}$ and $c_{j}$ is higher than the similarity between $r_{i}$ and $TR_{j}^{-}$, then $r_{i}$ is semantically closer to $c_{j}$ and should be traced to it. Conversely, if the similarity between $r_{i}$ and $TR_{j}^{-}$ is higher, then it should not be traced to $c_{j}$ as it is semantically closer to the negative examples. This procedure sets a different $\theta$ value for each $r_{i}$ based on randomly selected negative examples.
@@
(c) Maximum Delta Cutoff: In this method, we apply the following procedure. First, for each $r_{i}$, we sort the similarity values computed across the different provision $c_{j}\in\mathcal{C}$. Then, we compute delta values ($\Delta$) corresponding to the differences between each pair of consecutive similarity values and identify the largest $\Delta$ (i.e., the biggest gap in the computed similarities). To illustrate, consider the following example. Assume $r_{i}$ has similarity values of 0.98, 0.1, 0.3, and 0.7 with $c_{1}$, $c_{2}$, $c_{3}$, and $c_{4}$. We sort these values in descending order as follows: $c_{1}$: 0.98, $c_{4}$: 0.7, $c_{3}$: 0.3, $c_{2}$: 0.1. Next, we compute the $\Delta$ values: $\Delta(c_{1},c_{4})$=0.28, $\Delta(c_{4},c_{3})$=0.4, $\Delta(c_{3},c_{2})$=0.2. Based on these values, the largest $\Delta$ is 0.4 between $c_{3}$ and $c_{4}$. Finally, we set $\theta$ to the lower similarity value in the pair that yielded the largest $\Delta$. In the above example, we would set $\theta$ to 0.3 (the similarity value between $r_{i}$ and $c_{3}$). The largest $\Delta$ represents the most significant drop in similarity, indicating a potential boundary between relevant and irrelevant provision for $r_{i}$.
@@
The methods described above result in four variants of $\mathcal{K}$ashif, each determined by how $\theta$ is set. These variants are referred to as $\mathcal{K}$ashif constant, $\mathcal{K}$ashif dynamic, $\mathcal{K}$ashif Δ, and $\mathcal{K}$ashif tuned. We compare these variants in Section 4.

### v6

@@ 4 Lero Research Ireland centre for Software Research and University of Limerick, Ireland
Classifier or Prompt: A Case Study on Legal Requirements Traceability
@@ 1 Introduction
Technological advancements are significantly transforming software development across diverse domains, such as healthcare Caruana:15 . Software applications and automated assistants have become integral to our daily lives Zhan:22 . This evolution, driven by recent breakthroughs in artificial intelligence (AI), has led to increasing complexity in software systems ahmad2023requirements ; Feldt:18 . As technology advances, regulations are evolving in parallel to ensure that software systems are developed in accordance with ethical and legal standards. For example, the General Data Protection Regulation (GDPR) GDPR has been in effect since 2018 to address concerns regarding privacy and data protection. Although introduced by the European Union (EU), the GDPR has a global effect, affecting organizations (and software) outside the EU that process personal data of EU residents.
@@
Requirements Engineering (RE) plays a pivotal role in this landscape. RE is concerned with specifying and maintaining software requirements that outline the properties and functions of a system-to-be Pohl:11 . Legal compliance of software systems against applicable provisions can be addressed at different stages of software development. One scenario is to explicitly identify legal requirements early during the requirements elicitation phase, answering the question: “What legal obligations need to be satisfied by the system for it to be compliant?”. The elicited legal requirements can then be integrated into the software development process while maintaining traceability to the source legal provisions. As an alternative scenario, requirements engineers may need to verify compliance of existing software systems with legal provisions during the post-deployment stage, as new regulations have become applicable. In this case, they must answer the question “Does the system satisfy the regulation?”. To do so, engineers must analyze the regulation, identify the applicable legal provisions, and trace software requirements to those provisions. Both alternatives rely on requirements traceability analysis, an essential RE activity concerned with identifying and maintaining trace links between requirements and other artifacts throughout the software development lifecycle Meyer:22 . legal requirements traceability (LRT) is a special case where requirements are traced to provisions in a regulation and is the focus of this paper.
@@
To illustrate this concept, consider the following example. Consider a fictional mobility app named WeMobilize, which helps users book and share taxi rides. Originally a non-EU startup, WeMobilize is expanding into the EU and must therefore comply with the GDPR. This example is particularly relevant, as many businesses are globalizing and must adapt to data protection laws across jurisdictions. Fig. 1 illustrates how WeMobilize’s requirements (labeled REQ1–REQ5) can be traced to data protection policies under the GDPR GDPR . We identify trace links to provisions of the GDPR for REQ1 and REQ3—REQ5, visualized as dashed black lines. REQ2 has no trace link to GDPR in our example, as it does not involve processing users’ personal data.
@@
REQ1 involves collecting users’ personal information and must therefore be traced to two provisions: REG_DIR (for the direct collection of personal information) and REG_CON (for the explicit solicitation of users’ consent). Currently, consent is not part of REQ1, which prevents the identification of a trace link with REG_CON—a missing trace link is visualized as a red dashed line in the figure. Failing to identify this trace link may constitute a breach of the GDPR. Therefore, deploying WeMobilize as-is, without accounting for GDPR provisions, can lead to reputational and financial losses arising from GDPR violations. LRT can help identify potential non-compliance issues at an early stage, but it requires not only legal expertise but also substantial manual effort. Developing automated support is therefore beneficial to assist engineers and analysts in identifying applicable trace links.
@@
Requirements traceability is a well-explored problem in the RE literature, e.g., wang2018requirements ; tufail2017systematic . However, the extensive research on requirements traceability is not directly applicable to LRT due to the challenges outlined above. Despite the serious consequences of non-compliance, LRT has received limited attention from the community. Cleland-Huang et al. cleland:2010 ; Gibiec:2010 proposed a classifier that predicts trace links by computing the likelihood of requirements being traced to provisions based on indicator terms found in both provisions and requirements. Guo et al. Guo:17 focused on bridging the terminology gap between provisions and software requirements. They examined three methods, including the method by Cleland-Huang et al. mentioned above, as well as two others based on web mining and ontologies. The proposed methods aim to expand the terminology of the provisions by adding additional terms to better identify trace links.
@@
Existing traceability approaches for the LRT task, mentioned above, have several limitations. First, they leverage a statistical model based on TF-IDF (Term Frequency-Inverse Document Frequency), which primarily relies on lexical overlap between documents, where similarity is computed from the presence and frequency of overlapping words. This approach is mostly effective when the source and target texts share similar vocabulary. Still, they struggle to capture deeper semantic relationships, which are required in LRT, as the language used in legal texts often differs significantly from that in technical requirements. Second, existing techniques have yet to fully leverage the recent advances in natural language processing (NLP), evolving from latent semantic methods such as Latent Semantic Indexing (LSI) deerwester1990indexing and Latent Dirichlet Allocation (LDA) blei2003latent , which uncover hidden semantic structures through concept and topic modeling, toward more contextually rich approaches. We posit that (large) language models (LMs or LLMs) and contextual sentence embedding architectures offer capabilities beyond lexical or shallow syntactic matching, enabling them to capture nuanced legal and technical semantics crucial to LRT. While some recent work has applied language models to traceability tasks lin2021traceability , these efforts have not been tailored to the distinctive demands of aligning with legal requirements—such as interpreting domain-specific terminology, resolving cross-references, or bridging differences in stylistic and structural conventions between legal and technical texts. Architectures like Sentence Transformers can generate semantically rich representations that are better aligned with these needs but remain underutilized in this context. Third, prompt engineering has not been given sufficient attention; existing work often relies on overly simple prompts ronanki2024requirements ; hey2025requirements ; ge2025cross or on querying multiple LLMs fuchss2025beyond , which may not always be feasible in real-world deployments due to computational and cost constraints. Fourth, the evaluation relies on a single benchmark that may not fully capture the complexity of the legal domain in practice. To address these limitations, we propose two novel approaches in this paper, based on recent NLP technologies that utilize the Transformer architecture Vaswani:17 and LLMs. Similar to prior work, both approaches aim to predict trace links, and we evaluate their performance in a realistic scenario beyond the commonly used benchmark dataset.
@@
(3) We created a new dataset based on GDPR regulations. It includes 310 requirements covering diverse domains and requirements types. These requirements are traced to the GDPR, a complex regulation comprising 26 provisions on personal data protection, to which software requirements must comply.
(4) To further confirm $\mathcal{K}$ashif’s performance, we test it on our newly created unseen requirements documents. On this dataset, the base ST model (without additional fine-tuning) yields an average recall of 15%. In comparison, a pre-trained sentence transformer, with no exposure to the requirements traceability task, yields a nearly zero recall, as elaborated in Section 4.7. Driven by this observation, we propose our second solution, the final contribution of this paper, as explained next.
@@ 2 Background
Language Models (LMs). Language Modeling in NLP involves computationally determining the probability distribution of word sequences Jurafsky:20 . Given a sequence of words, an LM predicts the most likely next word, enabling it to generate text Alexandrescu:06 . For example, an LM would predict “Mat” as the most likely next word in the input sequence, “The cat sits on the [WORD]”. LMs are trained on large corpora of text to accurately estimate these probability distributions. State-of-the-art LMs are based on transformer architecture, which leverages self-attention mechanisms to weigh the significance of different parts of an input text relative to a given position Vaswani:17 . The attention mechanism determines which words in a sentence are more important in the context and assigns them greater attention. For instance, in the sentence “Mary, who used to live in Paris, loves wine.”, the attention is on Mary and wine. Building on transformer architectures, the Sentence Transformers framework (ST) Reimers:19 provides a set of pre-trained models that encode longer text sequences, such as sentences or paragraphs, into dense vector representations in a high-dimensional space. They produce contextual embeddings that capture the overall semantic essence of an entire input sequence.
More recently, generative LLMs have emerged as transformer-based language models that are significantly larger and trained on much more data. Examples on LLMs include OpenAI’s GPT (Generative Pre-trained Transformer) Radford:18 and LLaMa touvron2023llama ; touvron2023llama2 . These models can perform new tasks based on textual instructions (prompts) NEURIPS2020_1457c0d6 .
Machine learning (ML). Supervised learning is one of the most prominent paradigms in ML. In this paradigm, the ML algorithm is provided with labeled training data where each data point consists of an input vector (features) and the corresponding output label (or value). The algorithm learns patterns in the input features to make predictions from this training data. When trained on a sufficiently large dataset, the algorithm refines its predictions to classify the provided labels more accurately. Examples of ML classification algorithms include random forests, decision trees, support vector machines, and feedforward neural networks Goodfellow:16 .
@@ 3.2 $\mathcal{K}$ashif
Fig. 2 provides a comprehensive overview of the two phases comprising our approach. Phase A covers steps 1-3 and offers a developer’s perspective, focusing on building a traceability model for solving LRT. Step 1 prepares a training dataset of manually identified trace links. Step 2 selects a pre-trained model to customize for addressing LRT. Step 3 involves fine-tuning the LRT model. Phase B covers steps 4-6 and provides the perspective of an end user (e.g., a requirements analyst) assuming the availability of an LRT model. Step 4 preprocesses the input requirements document (RD). Step 5 applies the LRT model to compute the semantic similarity between each requirement in the RD and each provision. Step 6 predicts trace links. We explain these steps in detail next.
@@
In step 4, we preprocess the input requirements using a simple NLP pipeline composed of two modules: Tokenization and Sentence Splitting. The goal is to decompose the text into separate sentences. In our work, a requirement $r_{i}$ corresponds to a sentence generated by the NLP pipeline, which may or may not be grammatically correct. Using $\mathcal{K}$ashif to solve LRT for multi-sentence requirements is straightforward. A provision is traced to the requirement if it is traced to any sentence thereof. The intermediary output of this step is a set of $n$ requirements ($\mathcal{R}=\{r_{1},r_{2},\ldots,r_{n}\}$) from the input RD.
@@
(a) Constant Threshold: To predict a trace link, we use a predefined threshold, $\theta=0.5$. Specifically, a trace link is predicted if the similarity score exceeds $0.5$. This threshold is considered a reasonable rule of thumb, as evidenced by its previous application in the literature Yao2014 ; Corley2005 . Moreover, this threshold was chosen based on the fine-tuning approach used in $\mathcal{K}$ashif. During training, positive pairs are assigned a similarity score of 1, whereas negative pairs are assigned 0. The model is thus optimized to produce similarity scores closer to 1 for true links and closer to 0 for unrelated pairs. As a result, a threshold of 0.5 serves as the midpoint between these two extremes, making it a natural choice for distinguishing positive from negative links.
@@
(b) Dynamic Threshold: Another practical method to adjust $\theta$ involves curating a set of negative training examples, i.e., requirements that lack trace links. These requirements can be sourced from publicly available datasets or from different projects. However, for more accurate results, it is ideal to use requirements from the same project under analysis. Inspired by similarity-based classification proposed in the literature Amaral:21 , we select $\theta$ using the following procedure. For each provision $c_{j}\in\mathcal{C}$, we identify a set of negative training examples ($TR_{j}^{-}$), i.e., requirements $\{r_{1}^{\prime},\ldots,r_{k}^{\prime}\}$ that do not have trace links to $c_{j}$. We then compute the similarity between $r_{i}$ and $TR_{j}^{-}$ and set $\theta$ to the average cosine similarity between $r_{i}$ and $TR_{j}^{-}$. If the similarity between $r_{i}$ and $c_{j}$ is higher than the similarity between $r_{i}$ and $TR_{j}^{-}$, then $r_{i}$ is semantically closer to $c_{j}$ and should be traced to it. Conversely, if the similarity between $r_{i}$ and $TR_{j}^{-}$ is higher, then it should not be traced to $c_{j}$ as it is semantically closer to the negative examples. This procedure sets a different $\theta$ value for each $r_{i}$ based on randomly selected negative examples.
@@
(c) Maximum Delta Cutoff: In this method, we apply the following procedure. First, for each $r_{i}$, we sort the similarity values computed across the different provisions $c_{j}\in\mathcal{C}$. Then, we compute delta values ($\Delta$) corresponding to the differences between each pair of consecutive similarity values and identify the largest $\Delta$ (i.e., the biggest gap in the computed similarities). To illustrate, consider the following example. Assume $r_{i}$ has similarity values of 0.98, 0.1, 0.3, and 0.7 with $c_{1}$, $c_{2}$, $c_{3}$, and $c_{4}$. We sort these values in descending order as follows: $c_{1}$: 0.98, $c_{4}$: 0.7, $c_{3}$: 0.3, $c_{2}$: 0.1. Next, we compute the $\Delta$ values: $\Delta(c_{1},c_{4})$=0.28, $\Delta(c_{4},c_{3})$=0.4, $\Delta(c_{3},c_{2})$=0.2. Based on these values, the largest $\Delta$ is 0.4 between $c_{3}$ and $c_{4}$. Finally, we set $\theta$ to the lower similarity value in the pair that yielded the largest $\Delta$. In the above example, we would set $\theta$ to 0.3 (the similarity value between $r_{i}$ and $c_{3}$). The largest $\Delta$ represents the most significant drop in similarity, indicating a potential boundary between relevant and irrelevant provision for $r_{i}$.
@@
The methods described above yield four variants of $\mathcal{K}$ashif, each determined by the value of $\theta$. These variants are referred to as $\mathcal{K}$ashif constant, $\mathcal{K}$ashif dynamic, $\mathcal{K}$ashif Δ, and $\mathcal{K}$ashif tuned. We compare these variants in Section 4.